topfloorvamp:

Does anyone know how to convert degree measure into radians left in terms of pi?? (pls help a brotha ouut)

lowkey:

\[x~degree *\frac{ \pi }{ 180 }=radians\]

velmalovesshaggy145:

damn

velmalovesshaggy145:

lowkey is correct

lowkey:

For example, 90 degrees to radians: \[90~degrees * \frac{ \pi }{ 180 }=\frac{ 1 }{ 2 }\pi\]

lowkey:

which, a simple rememberance of the Unit Circle may come in handy for simple conversions as that

velmalovesshaggy145:

radians to degrees is #of radians times 180/pi

velmalovesshaggy145:

@lowkey wrote:
which, a simple rememberance of the Unit Circle may come in handy for simple conversions as that
Well said, also correct

topfloorvamp:

thank u for trying to help